This question is very important and I need solution for this issue with all the details a.b.c , and help me with all the details? BR/Ha

Question :

Consider an island with exclusive fishing rights to their own waters. The fishing industry is unregulated and the production function, representing the yearly catch, is given by f ( x )=22 x− x2 /40 where x is the number of boats launched. Yearly industry profit is split equally among fishing boats and the cost of launching a boat equals 80. Fish is sold at a world market price of p=40. A market failure is likely to occur in this market. Explain why and analyze the problem by comparing the profit maximizing solution for the fishing industry with “sustainable” fishing (revenue maximizing solution) and the free entry solution. Suggest a policy that leads to an efficient outcome.
Answer #1

free entry solution

at eqm, AR per boat = Cost of the boat

TR = P*Q

= 40*( 22x - x​​​​​​2​​​​​/40)

So AR = TR/x

= 40(22 - x/40)

So at eqm, 40(22- x/40) = 80

22 - x/40 = 2

20 = x/40

x* = 800 boats

Sustainable solution

at eqm . MR = MC

MR = dTR/dx

= 40(22 - 2x/40)

= 40*(22- x/20)

So at eqm , 40(22 - x/20) = 80

22 - x/20 = 2

x` = 400 boats

Thus free market solution leads to overuse of resources as the socially optimal level of boats employed for fishing is always lower than free market level, where exploitation of resources occur

policy that leads to efficient solution

1) enforce the property rights system

A common property like fishery should be privatised & property rights of usage should be assigned to the administration

Then it will stop the overuse of the resource

2) command & control system

Only the optimal number of boats should be allowed to enter in to sea to catch fish